Tilak made 44 and 72 on the first two tests in algebra, and has one test remaining. For Tilak to pass
the course, the average on the three tests must be at least 60. For what range of scores on his last test will Tilak pass the course?
Solve the problem by using an inequality.

Answer by venugopalramana(3286)

Travis made 44 and 72 on the first two tests in algebra and has one test remaining. The average on the three tests must be at least 60 for Travis to pass the course. Let s represent his score on the last test.
Write an inequality to describe each situation.
total score in 3 tests = 44+72+s = 116+s
average = (116+s)/3...this should be atleast 60..so
(116+s)/3 >=60
116+s>=180
s>=180-116=64
s>=64